Using python, Write a function max3 that takes three numbers as its arguments and returns the greatest of these three numbers. Your submitted file for this problem should include (a) function definition, obviously, and (b) the code that calls that function (for this problem you do not have to package that calling code into the main() function although you may if you want to), so that if I run your *.py file it computes and prints something (e.g. few test examples - you can hardcode a few function calls, no need to ask the user to enter the three numbers)
Answer #1
def max3(a,b,c):           #function definition for max3, it has three arguments a,b & c
    a = max(a,b,c)         #calculate the greater among three and store in a
    return a                #return the graetest number

x =max3(1,15,3)              #call the function by passing three argument and store the result in variable x
print(x)                      #print the result  

y = max3(10,20,30)
print(y)

z = max3(-1,-2,-3)
print(z)

Screenshot: